The average age of public jeepneys plying in Metro Manila is 15 years. Assume that the standard deviation is 8 years. If a random sample of 25 public jeepneys is chosen, Find the probability that the mean of jeepneys ages is between 12 and 19.


1
Expert's answer
2022-05-25T17:14:53-0400
"P(12<X<19)=P(Z<\dfrac{19-15}{8/\sqrt{25}})"

"-P(Z\le\dfrac{12-15}{8/\sqrt{25}})"

"=P(Z<2.5)-P(Z\le-1.875)"

"\approx0.99379-0.03034\approx0.9634"